All involved steps in the Cleveland Clinic case added up to 28 minutes per patient (assuming no waiting), and the bottleneck was the MRI machine with a processing time of 10 minutes. If the patients start to incur average wait times of 22 minutes due to fluctuations in arrival times and service times, on average how many patients would be in the clinic?
Answer #1

Bottleneck capacity = 60/10 =6 per hour ( which is production rate)

Flow time = waiting time + processing timw = 22+28 =50 minutes

Number of jobs in system = flow time x production rate

= 6 x (50/60) =5